OpenStudy (anonymous):

Show that if p is a prime number, and k is an integer\[ 2 \le k \le p\], then the sum of the products of each k-element subset of\[ \{1, 2, \ldots, p\}\] is divisible by p.

OpenStudy (kinggeorge):

Wouldn't this also be true for \(k=1\) since \(1+2+3+...+p≡0 \mod p\)? That is unless \(p=2\).

OpenStudy (kinggeorge):

How about rewriting \(\{1, 2, ..., p\}\) as \[S=\{1, 2, ..., \frac{p-1}{2}, -\frac{p-1}{2}, ..., -1, 0\}\]So any term in our sum of products that includes \(0\) is automatically divisible by \(p\). So we just need to show that for \(2\leq k\leq p-1\) the proposition holds for the set \(S\). I still can't tell if rewriting the set that way helps us or not though.

OpenStudy (anonymous):

Hint: Factor \[ x^p -x \text { in } Z_p \]

OpenStudy (kinggeorge):

Well let's see here. \(x^p-x=((x-1)(x-2)(x-3)...(x-(p-1))(x-p)\in\mathbb{Z}_p\) In this expression, we have every product of \(k\)-element subsets of \(\{1, 2, ..., p\}\) multiplied by some negative 1's.

OpenStudy (kinggeorge):

If we write this a different way, we get \[x^p-x^{p-1}(1+2+...+p)+x^2(1\cdot2+1\cdot3+...+(p-1)\cdot p)-x^{p-3}(...)+...-p!\]

OpenStudy (kinggeorge):

That should be \[x^p-x^{p-1}(1+2+...+p)+x^{p-2}(1\cdot2+1\cdot3+...+(p-1)\cdot p)-x^{p-3}(...)+...-p!\]

OpenStudy (kinggeorge):

If we set this equal to \(x^p-x\), we have that every term should be 0 mod p except for the last one with a factor of x. So every term is divisible by p except for the term that has the sum of products of \(p-1\) element subsets. In short, this proves it for \(2\leq k <p-1\) and \(k=p\). We're only missing \(k=p-1\).

OpenStudy (kinggeorge):

However, shouldn't that term be \(-1 \mod p\) by Wilson's Theorem? All the products in that sum have a factor of \(p\) except for the \((p-1)!\equiv -1\mod p\).

OpenStudy (kinggeorge):

I'm thinking that for \(k=p-1\), it should be \(-1\mod p\) since you have \[\frac{p!}{p}+\frac{p!}{p-1}+\frac{p!}{p-2}+...+\frac{p!}{2}+\frac{p!}{1} \mod p\]\[\equiv (p-1)!+p(\text{stuff}) \equiv -1+0\equiv-1\mod p\]

OpenStudy (kinggeorge):

For example. Let \(p=3\) and \(k=p-1=2\). Then we're summing \[1\cdot2+1\cdot3+2\cdot3=11\equiv 2\equiv-1 \mod p\]